A pet owner told the vet that he would drop his pet off between the hours of 9:30 AM and 1:45 PM. The vet plans to take lunch from 11:15 AM to 12:00 PM.

a. What is the probability that the owner will come during his lunch?

Round to 3 significant digits

b. When should he expect the owner to arrive?

  :  :  

(click to select)AMPM
Answer #1

This is a uniform distribution which ranges between 0 minute and 195 minutes. Hence,

a) P(Owner will come between 11:15 AM and 12:00 PM)

= 15 / (195 - 0)

= 15/195

= 0.077

b) Expected value = (0 + 195)/2 = 97.5

Hence,

Owner is expected to arrive 97.5 minutes after 9:30 i.e. 11 : 08 AM